A 5 kg mass is riding on a horizontal platform that is rotating with a period of 8 seconds. The mass is held in place by the friction at a distance of 5.20 meters from the center of rotation. The coefficient of friction between the mass and the platform must at least be?
1
Expert's answer
2019-12-12T06:21:53-0500

For the equilibrium:


"\mu mg=m\omega^2r"

"\mu =\frac{4\pi^2r}{gT^2}"

"\mu =\frac{4\pi^2(5.2)}{(9.8)(8)^2}=0.327"